OpenStudy (anonymous):

find the derivative y = ln(8x^3 + 3^8x)

OpenStudy (anonymous):

its absolute value

OpenStudy (anonymous):

ln absolute value

OpenStudy (anonymous):

\[y=\ln \left| 8x ^{3}+3^{8x} \right|\]

OpenStudy (anonymous):

how would you input this on wolphram.com

OpenStudy (anonymous):

unless one of you can solve this

OpenStudy (anonymous):

like this: ln|8x^3 + 3^8x|

OpenStudy (anonymous):

derive ln|8x^3 + 3^8x| for the derivative.

OpenStudy (anonymous):

unless one of you can solve this here

OpenStudy (zarkon):

note: if \[f(x)>0\] \[\frac{d}{dx}\ln(|f(x)|)=\frac{d}{dx}\ln(f(x))=\frac{f'(x)}{f(x)}\] if \[f(x)<0\] \[\frac{d}{dx}\ln(|f(x)|)=\frac{d}{dx}\ln(-f(x))=\frac{-f'(x)}{-f(x)}=\frac{f'(x)}{f(x)}\]

OpenStudy (anonymous):

forgot peranteces but its good thomas

myininaya (myininaya):

\[y=3^{8x}\] \[lny=\ln(3^{8x})\] \[lny=8xln3\] \[lny=8(\ln3)x\] \[\frac{y'}{y}=8\ln3\] \[y'=y*8\ln3\] \[y'=3^{8x}*8\ln3\]

OpenStudy (anonymous):

\[\frac{dy}{dx} = \frac {1}{8x^3 +3^{8x}}\times ( 24x^2 + \frac{d}{dx} 3^{8x} )\]

OpenStudy (anonymous):

then for the part second \[f = 3^{8x}\]\[\ln f = 8x \ln 3\]\[\frac{1}{f} \frac{df}{dx} = \ln3 \times 8\]then what you get \[\frac{df}{dx} = 3^{8x} \times \ln3\times 8\]Now all you need to do is to replace the Values .. : )

OpenStudy (anonymous):

\[\frac{dy}{dx} = \frac{1}{8x^2 + 3^{8x} }\times (24x^2 + 3^{8x}\times \ln3 \times 8)\]should be the derivative ... i'm doing it so long because I learned this fraction today and I like it ...lol :D